# Solvation Free Energies from Neural Thermodynamic Integration

### v1

@@ 4.1 Solute solvation of a Lennard-Jones fluid
We parameterize the interpolating potential in a way that enforces the boundary conditions at $t\in\{0,1\}$ (Máté & Fleuret, 2023).
$U_{t}(x)=b_{t}^{A}U_{\mathrm{solvent}}(x,a_{t}^{A})+b_{t}^{B}U_{\mathrm{solute}}(x,a_{t}^{B})+b^{AB}_{t}U_{\mathrm{solvent-solute}}(x,a^{AB}_{t})+b^{\mathrm{U}}_{t}U_{t}^{\theta}(x)$ (12)
where $U_{t}^{\theta}(x)$ is the same graph neural network as in Máté et al. (2024) with node features encoding the particle type (solute or solvent) of the nodes. The functions $a^{A}_{t},a^{B}_{t},a^{AB}_{t}$ denote softening parameters to respective components of the potential (Beutler et al., 1994). We use the following softened form of the LJ-potential
which corresponds to evaluating the Lennard-Jones potential at $\sqrt{r^{2}+a\sigma^{2}}$. When $a=0$, no softening is applied, and larger values of $a$ mean larger softening deformation of the original Lennard-Jones interactions. We enforce the boundary conditions of the the $b_{t}$ and $a_{t}$ functions by learning them with parametrization
$b_{t}^{A}=[1-t(1-t)]e^{\beta_{t}^{\mathrm{A}}}\quad b_{t}^{B}=[1-t(1-t)]e^{\beta_{t}^{\mathrm{B}}}\quad b_{t}^{\mathrm{AB}}=te^{t\beta_{t}^{\mathrm{AB}}}\quad b_{t}^{\mathrm{U}}=t(1-t)e^{\beta_{t}^{\mathrm{U}}}$ (14)
$a_{t}^{\mathrm{A}}=t(1-t)e^{\alpha_{t}^{\mathrm{A}}}\quad a_{t}^{\mathrm{B}}=t(1-t)e^{\alpha_{t}^{\mathrm{B}}}\quad a_{t}^{\mathrm{AB}}=(1-t)e^{t\alpha_{t}^{\mathrm{AB}}}$ (15)
where the $t\mapsto\alpha_{t},t\mapsto\beta_{t}$ functions are learned by a small neural network.
@@
The function we are aiming to learn is $U_{t}(x)$. The DSM objective controls its behavior with respect to variations of $x$, but there is no training signal controlling the behaviour of $U_{t}(x)$ with respect to changes to the temporal argument. To avoid large absolute values of $\partial_{t}U_{t}(x)$ we also include the $L_{2}$-norm of $\partial_{t}U_{t}$ to our optimization objective. Our training objective is thus
$\mathcal{L}(\theta)$ $=\mathbb{E}_{t\sim\mathcal{U}([0,1]),x_{t}\sim\rho_{t}}\Big{[}\mathcal{L}_{\mathrm{DSM}}+\kappa\mathcal{L}_{\mathrm{reg}}\Big{]}$ (16)
$\mathcal{L}_{\mathrm{DSM}}=\sigma_{t}^{2}||\nabla_{x}U_{t}(x_{t})-z||^{2}$ (17)
$\mathcal{L}_{\mathrm{reg}}=\frac{1}{2}(\partial_{t}U_{t}(x_{t}))^{2}$ (18)
@@
where $U_{t}$ depends on $\theta$ through Equation 12. Since $\partial_{t}U_{t}$ is the very thing we are interested in measuring, we use a small relative weight for this regularization in our experiments, $\kappa=10^{-10}$. The goal is not minimize the $L_{2}$-norm of $\partial_{t}U_{t}$, but to prevent it from diverging.
Interpolation
Given samples $x_{0},x_{1}\sim(\rho_{0},\rho_{1})$, we construct the function $I(x_{0},x_{1},t)$ in Equation (4) by translating both $x_{0}$ and $x_{1}$ to center the solute and then compute the optimal transport (OT) between the solvent particles with respect to the squared toroidal distance between the solvent particles. Note that we compute the OT-pairings between the solvent particles from two given snapshots $x_{0},x_{1}$ and not between datapoints in a batch as proposed by Tong et al. (2023) During the interpolation, the solvent particles move along the toroidal geodesics given by the OT-pairing , and we keep the solute frozen.
Figure 3: Solute solvation of a Lennard-Jones fluid. Left: Solvation free energy estimates by TI and neural TI as a function of number of reference simulations and training steps, respectively. Center: Detailed view at later training stages. Softening parameters of the LJ-interactions after training (right).
Results
@@
We compare our estimates with TI using evenly spaced time-slices along the time-interval $[0,1]$. We perform 51 simulations in total at $t\in\{0.00,0.02,0.04,...,1.00\}$, and find that the solvation free energy lies between $-21k_{B}T$ and $-19k_{B}T$. The rolling estimate of neural TI converges to $\approx-19.5k_{B}T$. The time evolution of these values are shown in the left subplot of Figure 3. The right subplot of Figure 3 displays the learned softening schedule of the solvent-solvent and solute-solvent Lennard-Jones interactions. The learned softening functions show that our interpolation does not correspond to the usual linear interpolation $(1-t)U_{0}+tU_{1}$, even the interactions within the solvent are softened and decoupled along the trajectory.
4.2 Hydration free energies
To move to more realistic chemical systems, we now extend the experimental setup to compute hydration free energies of small molecules, namely water and methane, from atomistic simulations. We work with rigid molecules, i.e., we consider Lennard-Jones and Coulomb, but not bonded, interactions. Microstates of single molecules are thus described by the the position of their central atom (oxygen for water, carbon for methane) and their orientation, i.e., a group element of $SO(3)$.
System description
The solvent consists of 216 TIP4P water molecules (Jorgensen et al., 1983) at density 1 g/cm3, resulting in a simulation box of size $18.62{\AA}\times 18.62{\AA}\times 18.62{\AA}$. The hydration of water—solvation of water in water—simply consists of coupling one extra TIP4P molecule in the box. For methane, we rely on the atomistic OPLS (OPLS-AA) force field (MacKerell Jr et al., 1998). Coulomb interactions are modelled with the reaction field method (Tironi et al., 1995)
$E_{\text{Coulomb}}(r)$ =$=\begin{cases}k_{C}{q_{1}q_{2}}\left(\frac{1}{r}+k_{\text{rf}}r^{2}-c_{\text{rf}}\right),\quad\text{if}\quad r\leq r_{\text{cutoff}}\\ 0,\quad\text{if}\quad r>r_{\text{cutoff}}\end{cases}$ (19)
$k_{\text{rf}}$ $=\left(\frac{1}{r_{\text{cutoff}}^{3}}\right)\left(\frac{\epsilon_{\text{water}}-1}{2\epsilon_{\text{water}}+1}\right)$ (20)
$c_{\text{rf}}$ $=\left(\frac{1}{r_{\text{cutoff}}}\right)\left(\frac{3\epsilon_{\text{water}}}{2\epsilon_{\text{water}}+1}\right)$ (21)
where $r_{\text{cutoff}}$ is the cutoff radius, which we set to $8.5\AA$, $k_{C}$ is Coulomb’s constant, $q_{1},q_{2}$ are the charges of the atoms, $r$ is the distance between them and $\epsilon_{\text{water}}$ is the dielectric constant of water, $\epsilon_{\text{water}}=78$. To soften this potential with softening parameter $a$, we define
$E^{\text{soft}}_{\text{Coulomb}}(r,a)=E_{\text{Coulomb}}(\sqrt{r^{2}+a})$ (22)
To learn the interpolating potential we use the ansatz of equation (12), which now contains both the soft Coulomb interactions defined by equation (22) in addition to the soft Lennard-Jones interactions defined by equation (13). Importantly, unlike the usual practice of TI, we do not prescribe to couple the Lennard-Jones interactions first, and the Coulombic interactions later, we let the network learn the complete coupling schedule with separate softening parameters for the two interactions.
Interpolation
Each molecule is represented by the position of its central atom (oxygen for water, carbon for methane) in $\mathbb{T}^{3}$ and its orientation in $SO(3)$. To interpolate between two snapshots, $x_{0}$ and $x_{1}$, we perform the same steps as in the previous section for the central atoms, and use the geodesic interpolation on $SO(3)$ between the orientations. To further minimize the rotational transport cost, we use the fact that both the rigid water and methane molecules have non-trivial symmetric groups, $\mathbb{Z}_{2}$ and $S_{4}$, respectively. Applying any of these transformation does not change the physical system itself, only its numerical representation [^1] Geometrically speaking, this means that the configuration space of a single rigid molecule can be reduced from $\mathbb{T}^{3}\times SO(3)$ to $\mathbb{T}^{3}\times SO(3)/G$, where $G\in\{\mathbb{Z}_{2},S4\}$ is the inner symmetry group of the molecule. thus we compute the geodesic distances corresponding to the physically equivalent orientations and choose the one with the minimal geodesic distance.
Architecture
@@
We benchmark our calculations on previously reported calculations of the hydration free energy for both water and methane: $-10.7k_{B}T$ (Jorgensen et al., 1989) and $3.4k_{B}T$ (Mobley et al., 2009), respectively. Figure 4 (right panel) compares our rolling neural TI estimates (equation (6)) with the reference values mentioned. For each system we provide two individual calculations with different initial random seeds. For the moving average we use a window size of 4,000 batches of batch size 8, i.e., 32,000 samples. The runs for water and methane use the same architecture of the neural network potential, same configuration, the runs only differ by the solute. The runs for the water and methane visibly diverge after 5,000 traing steps and stay in the ±$\pm 4k_{B}T$ range of the experimental value after 10,000 training steps. The estimates for water appear to be more consistent, this is likely due to the fact that the water-in-water system is more homogeneous than the methane-in-water system. All runs were performed on an Nvidia 3090 and took about 4 hours to reach the plotted 50,000 training steps.
@@
Figure 4: Illustration of the coupled water (left) and methane (center) molecules. The oxygen of the solute water molecule is colored blue to distinguish it from the solvent. Hydration free energy estimates of water and methane in TIP4P water (right). The dashed lines denote the experimental hydration free energies and the the curves of the same our rolling estimate from 2 different random seeds for each solute.
5 Conclusion
@@
We build on the ideas introduced by Máté et al. (2024) to estimate free energy differences between arbitrary pairs of target distributions. This means that diffusion models are replaced by their generalization, stochastic interpolants (Albergo & Vanden-Eijnden, 2023), a natural framework for our purposes. The method now relies on two, possibly non-trivial, Hamiltonians over the same configuration space, samples from their Boltzmann distributions, and an interpolating procedure between samples from these two distributions. The upside of this approach, compared to traditional TI, is that we only need data at the end-point Boltzmann distributions, entirely alleviating the need for intermediate reference MC or MD simulations.
Additionally, we streamlined the process by incorporating techniques like the rolling estimate of free energy differences (equation 6) and the regularization of the temporal derivative (equation 16), both of which we found beneficial during our experiments. In §3 we provide some theoretical insight into the characteristics of the learning paradigm.
The method was verified by accurately estimating the solvation free energy of a Lennard-Jones solute in a Lennard-Jones solvent and the hydration free energies of water and methane. We restricted our attention to rigid molecules, future work could explore how to include bonded interactions into the approach.
@@ Acknowledgments
We thank Aleksander Durumeric, Daniel Nagel and Youssef Saied for discussions. BM acknowledges financial support by the Swiss National Science Foundation under grant number CR - SII5 - 193716 (RODEM). TB acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).

### v2

@@ III Method
We parameterize the interpolating potential in a way that satisfies the boundary conditions at $t\in\{0,1\}$ [40].
$U_{t}(x)=\,b_{t}^{A}U_{\mathrm{solvent}}(x,a_{t}^{A})+b_{t}^{B}U_{\mathrm{solute}}(x,a_{t}^{B})+b^{AB}_{t}U_{\mathrm{solvent-solute}}(x,a^{AB}_{t})+b_{t}U_{t}^{\theta}(x),$ (20)
where $U_{t}^{\theta}(x)$ is a neural network with trainable parameters. We enforce the boundary conditions of the softening $t\rightarrow a_{t}^{A},a^{AB}_{t},a_{t}^{B}$ and coupling functions $t\rightarrow b_{t}^{A},b^{AB}_{t},b_{t}^{B}$ by the parametrization
$b_{t}^{A}=[1-t(1-t)]e^{\beta_{t}^{\mathrm{A}}}$ $\quad b_{t}^{B}=[1-t(1-t)]e^{\beta_{t}^{\mathrm{B}}}$ (21)
$b_{t}^{\mathrm{AB}}=te^{t\beta_{t}^{\mathrm{AB}}}$ $\quad b_{t}=t(1-t)e^{\beta_{t}}$ (22)
$a_{t}^{\mathrm{A}}=t(1-t)e^{\alpha_{t}^{\mathrm{A}}}$ $\quad a_{t}^{\mathrm{B}}=t(1-t)e^{\alpha_{t}^{\mathrm{B}}}$ (23)
$a_{t}^{\mathrm{AB}}$ $=(1-t)e^{t\alpha_{t}^{\mathrm{AB}}},$ (24)
where the $t\mapsto\alpha_{t}^{*},t\mapsto\beta_{t}^{*}$ functions are learned by a small multilayer perceptron. These functional forms ensure that at $t=0$ and $t=1$ the solute and solvent interactions are fully coupled and no softening is applied to them, while the solvent–solute interactions are fully decoupled at $t=0$ and coupled at $t=1$. We predict such a set of parameters for each interaction separately.
Interpolation
We use the term constituent to refer to the solvent LJ particles in the case of the LJ solvent and to the rigid water molecules in the case of the water solvent. Given samples $x_{0},x_{1}\sim(\rho_{0},\rho_{1})$, we construct the function $I(x_{0},x_{1},t)$ in Equation (4) by translating both $x_{0}$ and $x_{1}$ to center the solute and pair the solvent constituents using on the optimal pairing with respect to the squared toroidal distance between the snapshots $x_{0},x_{1}$. In the case of water solvent, we consider the distances between the oxygen atoms. Note that we compute the OT-pairings between the solvent constituents from two given snapshots $x_{0},x_{1}$ and not between datapoints in a batch as proposed by [41]. We keep the solute frozen throughout the interpolation at its state in $x_{1}$. During the interpolation, the positions of the solvent constituents move along the toroidal geodesics given by the OT-pairing. In case of the water solvent, the orientations of the water molecules are also interpolated using the geodesic interpolation on $SO(3)$. To further minimize the rotational transport cost we exploit the inner $\mathbb{Z}_{2}$-symmetry of the rigid water molecule. Applying this flip transformation does not change the physical system itself, only its numerical representation, thus we compute the geodesic distances corresponding to the physically equivalent orientations and choose the one with the minimal geodesic distance. Geometrically speaking, this means that the configuration space of a single rigid water molecule can be reduced from $\mathbb{T}^{3}\times SO(3)$ to $\mathbb{T}^{3}\times SO(3)/\mathbb{Z}_{2}$.
Architecture
For the LJ system, we parametrize $U_{t}^{\theta}$ in equation (20) with the same graph neural network as in [24] with node features encoding the particle type (solute or solvent) of the nodes. In the hydration experiment, to push the method further, we write $U_{t}^{\theta}$ as a sum of pairwise interactions between all pairs of atoms in the system. First, we label each atom into one of the following four categories: (1) central atom of solute; (2) hydrogen in solute; (3) hydrogen in solvent; (4) oxygen in solvent. Then we consider the edge between all pairs of atoms, excluding pairs within the same molecule and label these edges into seven partitions, $\{P_{1},...,P_{7}$}, by the labels of their endpoints. We then predict an energy for each pair by using seven different multi-layer perceptrons, one for each category, taking as input the distance between the edge endpoints, the distance from the solute, and time as input,
$U_{t}^{\theta}=\sum_{k}\sum_{(i,j)\in P_{k}}f_{\psi_{k}}(t,d_{ij},d_{i},d_{j}),$ (25)
where $k$ runs over the seven partitions of the edges, $t$ is time, $d_{ij}$ denotes the distance between atoms $i$ and $j$ and $d_{i}$ the distance between atom $i$ and the central atom of the solute. The functions $f_{\psi_{k}}$ with parameters $\psi_{k}$ are parametrized by multilayer perceptrons of with hidden layers $[96,96]$.
@@
The function we are aiming to learn is $U_{t}(x)$. The SM objective controls its behavior with respect to variations of $x$, but there is no training signal controlling the behavior of $U_{t}(x)$ with respect to changes to the temporal argument. To avoid large absolute values of $\partial_{t}U_{t}(x)$ we also include the $L_{2}$-norm of $\partial_{t}U_{t}$ to our optimization objective. Our training objective is thus
$\mathcal{L}(\theta)$ $=\mathbb{E}_{t\sim\mathcal{U}([0,1]),x_{t}\sim\rho_{t}}\Big{[}\mathcal{L}_{\mathrm{SM}}+\kappa(\partial_{t}U_{t}(x_{t}))^{2}\Big{]},$ (26)
@@
where $U_{t}$ depends on $\theta$ through equation 20 and $\mathcal{L}_{\mathrm{SM}}\in\{\mathcal{L}_{\mathrm{DSM}},\mathcal{L}_{\mathrm{TSM}}\}$. Since $\partial_{t}U_{t}$ is the very thing we are interested in measuring, we use a small relative weight for its regularization in our experiments, $\kappa=10^{-10}$. The goal is not to minimize the $L_{2}$-norm of $\partial_{t}U_{t}$, but rather to prevent it from diverging.
IV Results
LJ solute in an LJ fluid
We consider a system of 512 particles interacting via an LJ potential. The LJ parameters of the solvent are denoted by $(\varepsilon_{A},\sigma_{A})$ while for the solute $(\varepsilon_{B},\sigma_{B})=(2\varepsilon_{A},2\sigma_{A})$. The simulation box is of size $9\sigma_{A}\times 9\sigma_{A}\times 9\sigma_{A}$ with periodic boundary conditions. To train the model we employ denoising score matching.
@@ IV Results
We compare our estimates with TI using evenly spaced time-slices along the time-interval $[0,1]$. We perform 51 simulations in total at $t\in\{0.00,0.02,0.04,...,1.00\}$, and find that the solvation free energy lies between $-21k_{B}T$ and $-19k_{B}T$. The rolling estimate of neural TI converges to approximately $-19.5k_{B}T$. The time evolution of these values are shown in the left subplot of Figure 3. The solvation free energy estimate uses data only from the two endpoints and shows significantly lower variance than the associated standard TI computation, that requires between 5 and 50 interpolating timesteps. The right subplot of Figure 3 displays the learned softening schedule of the solvent–solvent and solute–solvent Lennard-Jones interactions. The learned softening functions show that our interpolation does not correspond to the usual linear interpolation $(1-t)U_{0}+tU_{1}$, even the interactions within the solvent are softened and decoupled along the trajectory.
Figure 3: Solute solvation of a Lennard-Jones fluid. Left: Solvation free-energy estimates by TI and neural TI as a function of number of reference simulations and training steps, respectively. Center: Detailed view at later training stages. Softening parameters of the LJ-interactions after training (right).
Hydration free energies
@@
We benchmark our calculations on previously reported calculations of the hydration free energy for both water and methane: $-10.7k_{B}T$ [44] and $3.4k_{B}T$ [45], respectively. Figure 4 (right panel) compares our rolling neural TI estimates (equation (12)) with the reference values mentioned. For each system we provide 3 individual calculations with different initial random seeds. For the moving average we use a window size of 2,000 batches of batch size 24, i.e., 48,000 samples. The runs for water and methane use the same architecture of the neural network potential, same configuration, the runs only differ by the solute. The runs for the water and methane visibly diverge after 2,000 training steps and the means over the 3 respective seeds stay in the ±$\pm 2k_{B}T$ range of the experimental value after 10,000 training steps. The standard error of the mean for both systems is below $3k_{B}T$. All runs were performed on an Nvidia 3090 and took approximately three hours to reach the reported 24,000 training steps.
@@
Figure 4: Illustration of the coupled water (left) and methane (center) molecules. Hydration free-energy estimates of water and methane in TIP4P water (right). The dashed lines denote the experimental hydration free energies and the the curves of the same our rolling estimate from three different random seeds for each solute.
Given that the two solutes are of comparable size and complexity, we may assume that two systems, and in particular their solute–solvent coupling process, share some similarities. In the ideal case, a model trained on one of the solutes could give a reasonable prediction when evaluated on the other. We test this with the trained models and confirm that this is indeed the case: both models learn features that are useful when modeling the coupling of the other solute and predict the hydration free energy of the unseen solute within a $4k_{B}T$ accuracy. We attribute this transferability to our ansatz (equation 20), which enforces the target Hamiltonians, along with the contribution of solvent-solvent interactions. The learned softening of the endpoints and the accurate modeling of solvent-solvent interactions throughout the coupling process are likely key factors underlying the observed transferability.
All numerical values are reported in Table 1.
Table 1: Experimental and neuralTI-estimated hydration free energies. All the neuralTI estimates are computed over $64,000$ $(t,x_{t})$ samples from $(\mathcal{U}([0,1]),\rho_{t})$ and averaged over three seeds. The colored cells contain the transfer predictions, i.e., predictions on solutes that the model was not trained on.
$\Delta F_{\text{hydration}}^{\text{water}}$ $\Delta F_{\text{hydration}}^{\text{methane}}$
experimental reference value $-10.7k_{B}T$ $3.4k_{B}T$
neural TI with a model trained on water $-11.5k_{B}T$cellcolorred!20 $1.6k_{B}T$
[…]
V Conclusion
@@
We build on ideas introduced by Máté et al. [24] to estimate free-energy differences between arbitrary pairs of target potentials by performing thermodynamic integration with a neural-network potential. Motivated by generalizations of diffusion models [25, 26], our method relies on two, possibly non-trivial, Hamiltonians over the same configuration space, samples from their Boltzmann distributions, and an interpolating procedure between samples from these two distributions. The upside of this approach, compared to traditional TI, is that we only need reference Boltzmann distributions for the end points, entirely alleviating the need for intermediate reference MC or MD simulations. Additionally, we simplify and improve the neuralTI approach by incorporating techniques like the rolling estimate of free-energy differences (equation 12), target score matching, and the regularization of the temporal derivative (equation 26), each of which we found beneficial during our experiments.
The method is verified by accurately estimating the solvation free energy of a Lennard-Jones solute in a Lennard-Jones solvent using only data from the endpoint simulations, matching the accuracy of standard TI, which typically requires tens of intermediate simulations. We also predict the hydration free energies of rigid water and methane, showing good agreement with experimental values, using a simple three-body potential. Future work could explore introducing bonded interactions into the system, which would eliminate the need for a non-Euclidean configuration space but would require handling forces acting on different length scales. Finally, we observed promising transferability of the learned potential between these two solutes.
@@ Acknowledgments
We thank Aleksander Durumeric, Daniel Nagel, Lorenz Richter, and Youssef Saied for discussions. BM acknowledges financial support by the Swiss National Science Foundation under grant number CR - SII5 - 193716 (RODEM). TB acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster).